A submarine left Hawaii travelling west four hours before an aircraft carrier. The aircraft carrier traveled in the opposite direction going 8 mph slower than the submarine for three hours which time the ships were 206 miles apart. Find the submarine's speed.
Answer by ikleyn(52794)

.
4x + 3*(x + (x-8)) = 206       <<<---===  the sum of the two distances


4x + 3x + 3x - 24 = 206


10x = 206 + 24


10x = 230


x = 230/10 = 23 miles per hour is the submarine' speed.     ANSWER